OpenStudy (anonymous):

I need help with this one, too. -9(b+3)+28b-7)=0

OpenStudy (anonymous):

-9b-27 +28b-7=0 19b-34=0 19b=34 b=34/19

OpenStudy (anonymous):

must let me know if any doubt

OpenStudy (anonymous):

I got that,too. I'm just checking my work. Can you help me with my other problem that I posted?

OpenStudy (anonymous):

ok.. that i also solved ..plz check

OpenStudy (anonymous):

is there any thing else